Please help me write an equation for this problem and then solve. You and your family are taking a trip. You begin with 17 gallons of gasoline in the fuel tank. Your car uses 3 gallons of gasoline per hour of driving. You will stop to refuel when there is exactly 2 gallons of gasoline remaining in the tank. How many hours can you travel before you stop to refuel?
Answer by rfer(16322)

(17-2)/3=
15/3=5 hr